# Post-adiabatic waveform-generation framework for asymmetric precessing binaries

### v2

@@ I Introduction
Gravitational self-force (SF) theory originally grew out of the need to model extreme-mass-ratio inspirals (EMRIs) for the planned gravitational-wave detector LISA [1]. After decades of progress [2], it has now matured into a practical framework for building fast, accurate gravitational waveform models [3, 4, 5, 6]. This development is exemplified by the FastEMRIWaveforms (FEW) software package, which can generate LISA-length EMRI waveforms on a timescale of $\sim 10$ms [7, 3, 8, 9, 10]. Moreover, while it was traditionally motivated by EMRIs, with mass ratios as extreme as $\sim 1:10^{7}$ [11], SF theory exhibits good agreement with fully nonlinear numerical relativity simulations even at mass ratios $\sim 1:10$ [12, 13, 14, 15, 16, 5, 17, 18, 19, 20, 21], and it is now recognized as one of the most viable methods of directly modelling intermediate-mass-ratio inspirals [22] as well as an important tool for calibrating effective models that cover the entire binary parameter space [23, 24, 25].
@@
It is also widely expected that meeting the accuracy requirements of future gravitational-wave detectors [33, 34, 35, 36, 37, 38, 39, 40, 41] will require carrying the SF expansion to second perturbative order in the mass ratio [22, 42], necessitating second-order calculations for generic configurations on a Kerr background. Moreover, at second order in SF theory, the spin of the secondary object enters [43] and contributes significantly to the waveform phase [44]. Over the past several years, this importance of the secondary spin has spurred rapid progress toward SF waveform models for spinning secondaries on generic orbits around spinning primaries [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 44, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73]. Our aim in this paper is to describe a complete waveform-generation framework for that generic scenario, focusing on the impact of the secondary’s spin.
@@ I.1 Multiscale waveform generation with a nonspinning secondary
[…]
@@
where ˙$\dot{\cal E}^{(1)}_{\cal H}$ and ˙$\dot{\cal L}^{(1)}_{\cal H}$ are the leading-order fluxes of energy and angular momentum into the primary’s horizon. We use rings over variables that cleanly separate slow evolution from rapid oscillations, avoiding oscillations in the mode amplitudes $\mathring{h}^{(n)}_{\bm{k}}$ and in the right-hand side of Eqs. (2) and (3).
In practical models, all of the inputs, $\mathring{h}^{(n)}_{\bm{k}}$, $F^{(n)}_{i}$, ˙$\dot{\cal E}^{(n)}_{\cal H}$, and ˙$\dot{\cal L}^{(n)}_{\cal H}$, are pre-computed on a grid of $J_{i}$ values. [^1] The background spin must also be included as an axis on this parameter-space grid, while the background mass only provides an overall scale. The corrections $\delta m_{1}$ and $\delta\chi_{1}$ do not need to be included as axes because $h^{(2)}_{\bm{k}}$ and $F^{(1)}_{i}$ are linear in these corrections, meaning one can pre-compute coefficients of $\delta m_{1}$ and $\delta\chi_{1}$ as functions of $J_{i}$ and of the background spin. Waveforms are then rapidly generated by solving Eqs. (2)–(5) and summing the modes in Eq. (1). Equations (2)–(5) can be solved rapidly because all dependence on the phases $\mathring{\psi}^{i}$ has been eliminated through the choice of variables $(\mathring{\psi}^{i},\mathring{\pi}_{i})$, avoiding the need to resolve oscillations on the orbital time scales $\sim 2\pi/\Omega^{i}$ [79].
In Eqs. (2)–(5), numeric labels in parentheses denote the post-adiabatic (PA) order at which the term enters the evolution of the orbital phases [74, 76], and $F^{(n)}_{i}$ are referred to as $n$PA forcing functions. To understand this order counting, note Eqs. (2)–(5) imply that the orbital phases, and hence the waveform phase [21], have an asymptotic expansion of the form
@@
Although this expansion has limited accuracy compared to a direct solution of Eqs. (2)–(5[5], it provides a simple estimate of the impact of any given term in the dynamics: $n$PA forcing functions directly contribute to the $n$PA phase $\mathring{\psi}_{(n)}^{i}$.
@@ II.1 Field equations and effective metric
[…]
@@
Importantly, the stress-energy terms (10a) and (10b) take the form of a spinning particle in a certain effective vacuum metric $\hat{g}_{\alpha\beta}$ rather than in the external background $g_{\alpha\beta}$. The proper time $\hat{\tau}$, four-velocity $\hat{u}^{\alpha}\equiv\frac{dz^{\alpha}}{d\hat{\tau}}$, metric determinant $\hat{g}$, and covariant derivative ^$\hat{\nabla}_{\alpha}$ are all defined from $\hat{g}_{\alpha\beta}$. The effective metric itself is defined by subtraction of suitably defined, singular self-fields $h^{{\rm S}(n)}_{\alpha\beta}$ from the physical metric. Given the physical metric from Eq. (8), which we can write as
@@ IV.1 Orbital motion
The analysis in Ref. [76] is valid for generic accelerated orbits in Kerr, and we will make extensive use of equations in it. However, our treatment differs from Ref. [76] in that we work with $t$ as our time parameter from the beginning, while Ref. [76] begins with a complete analysis in terms of Mino time $\lambda$ and then transforms to variables based on $t$ [^5] We also make several changes of notation from Ref. [76]: $p^{i}\to\pi_{i}$, $\varphi_{i}\to\mathring{\psi}^{i}$, $p^{i}_{\varphi}\to\mathring{\pi}_{i}$, and $\mathscr{P}^{\alpha}\to\varpi_{I}$, along with various less noteworthy alterations. Differential equations in terms of the two variables are related by the factor
@@ IV.1.1 Osculating geodesic equations
$\cos\theta(\psi^{i},\pi_{i})$ $=\cos\iota\cos\psi^{\theta},$ (81b)
@@
[…]
@@ IV.1.2 Averaging transformation
$F^{(0)}_{i}+\Omega^{j}_{(0)}\frac{\partial\delta\pi_{i}}{\partial\mathring{\psi}^{j}}+\Omega_{s(0)}\frac{\partial\delta\pi_{i}}{\partial\mathring{\psi}_{s}}$ $=f^{(0)}_{i}(\psi^{j}_{(0)},\mathring{\varpi}_{j},\tilde{\psi}_{s(0)}),$ (97b)
@@
[…]
@@
Note that the criterion of removing oscillations only determines the oscillatory parts of the functions $\Delta\psi^{i}$, $\delta\pi_{i}$, and $\Delta\psi_{s}$ (and their higher-order analogues); their averages $\langle\Delta\psi^{i}\rangle$, $\langle\Delta\pi_{i}\rangle$, and $\langle\Delta\psi_{s}\rangle$ do not enter into equations such as (102). Hence, these non-oscillatory pieces of the transformation can be chosen arbitrarily. However, such a choice affects the subleading terms in the equations of motion (87), (88), and (92). Concretely, we see from Eqs. (103) that the non-oscillatory functions have the following contribution to the subleading frequency corrections and forcing functions:
@@ IV.2 Field equations
$\left(\frac{\partial}{\partial t}\right)_{\!\bm{k},q}$ $=-i\bigl{(}k_{i}\Omega^{i}_{(0)}+q\,\Omega_{s(0)}\bigr{)}$
@@
$T^{(m)}_{\alpha\beta}(\mathring{\psi}^{i},\mathring{\varpi}_{I},\mathring{\psi}_{s},x^{i},\varepsilon)=m_{2}\frac{\hat{g}_{\alpha\mu}\hat{g}_{\alpha\nu}\dot{z}^{\mu}\dot{z}^{\nu}}{\sqrt{-\hat{g}_{\rho\sigma}\dot{z}^{\rho}\dot{z}^{\sigma}}}\frac{\delta^{3}(x^{i}-z^{i})}{\sqrt{-\hat{g}}}$ (121)
@@
$\cos\theta_{(0)}$ $=\cos\mathring{\iota}\cos\psi^{\theta}_{(0)},$ (126b)
@@
Given the expansions of the Einstein tensor and stress-energy tensor, the Einstein equation (14) divides into first- and second-order equations for the Fourier mode coefficients $\mathring{h}^{(1,\bm{k})}_{\alpha\beta}$ and $\mathring{h}^{(2,\bm{k},q)}_{\alpha\beta}$:
@@
Equations (140) and (141) can be solved on a grid of $\mathring{\pi}_{i}$ (and background spin) values. From the solutions $\mathring{h}^{(1,\bm{k})}_{\alpha\beta}$ and $\mathring{h}^{(2,\bm{k},q)}_{\alpha\beta}$, the forcing functions $F^{(n)}_{i}$ and frequency corrections $\Omega^{i}_{(1)}$ can be computed, and the waveform amplitudes can be read off the asymptotic values of $\mathring{h}^{(1,\bm{k})}_{\alpha\beta}$ and $\mathring{h}^{(2,\bm{k},q)}_{\alpha\beta}$. Using this precomputed data, Eqs. (95) can be solved to evolve the system through parameter space and generate a waveform, as described in the Introduction. We return to the waveform generation scheme in Sec. V.
@@ IV.3 Linear-in-spin effects
where $h^{(2\text{-}\chi_{2})}_{\alpha\beta}$ is the metric perturbation sourced by the spin. The first source term on the right-hand side is confined to the libration region containing the particle’s orbit, while the second source term is distributed over the entire spacetime. As highlighted in Ref. [44] and discussed in the next section, there is considerable advantage in choosing a (phase-space) gauge that eliminates this non-compact term.
@@ IV.4.1 Fixed frequencies and fixed turning points
    $\mathring{\psi}^{i}$ vanishes at turning points (or what would be turning points if $\mathring{\pi}_{i}$ were fixed). More generally, the phases $\mathring{\psi}^{i}$ vanish at the same locations as $\psi^{i}$.
@@
    either $\mathring{\pi}_{i}$ is geodesically related to $\Omega_{s}\equiv\frac{d\mathring{\psi}_{s}}{dt}$, making $\Omega_{s}=\Omega_{s(0)}(\mathring{\pi}_{i})$, or $\mathring{\psi}_{s}$ vanishes where $\tilde{\psi}_{s}$ vanishes.
We stress that our condition of ‘fixed turning points’ is unrelated to the condition with the same name in Ref. [60]; ours is a condition on the values of the radial and polar phases $\mathring{\psi}^{r}$ and $\mathring{\psi}^{\theta}$ at turning points, while Ref. [60]’s is a condition on the values of Boyer-Lindquist radius $r$ and polar angle $\theta$ at turning points.
@@ V.1 Waveform generation scheme
$h$ $=\sum_{\bm{k}\in\mathbb{Z}^{3}}\Bigl{[}\varepsilon\mathring{h}^{(1)}_{\bm{k}}(\mathring{\pi}_{i},\theta,\phi)+\varepsilon^{2}\mathring{h}^{(2\text{-}\not\chi_{2})}_{\bm{k}}(\mathring{\varpi}_{i},\theta,\phi)$
@@
with $\mathring{h}^{(1)}_{\bm{k}}=\lim_{r\to\infty}\Bigl{(}rh^{(1,\bm{k})}_{\alpha\beta}\bar{m}^{\alpha}\bar{m}^{\beta}\Bigr{)}$, for example. Equation (179) extends Eq. (1) to include the secondary spin. Here $\mathring{h}^{(2\text{-}\not\chi_{2})}_{\alpha\beta}$ denotes the $\chi_{2}$-independent part of the second-order metric perturbation, identical to the second-order term in Eq. (1). We have also divided the linear spin contribution—the solution to Eq. (152)—into pieces proportional to $\chi_{\parallel}$ and $\chi_{\perp}$, respectively.

### v3

@@ I Introduction
Gravitational self-force (SF) theory originally grew out of the need to model extreme-mass-ratio inspirals (EMRIs) for the planned gravitational-wave detector LISA [1]. After decades of progress [2], it has now matured into a practical framework for building fast, accurate gravitational waveform models [3, 4, 5, 6]. This development is exemplified by the FastEMRIWaveforms (FEW) software package, which can generate LISA-length EMRI waveforms on a timescale of $\sim 10$ms [7, 3, 8, 9, 10, 11]. Moreover, while it was traditionally motivated by EMRIs, with mass ratios as extreme as $\sim 1:10^{7}$ [12], SF theory exhibits good agreement with fully nonlinear numerical relativity simulations even at mass ratios $\sim 1:10$ [13, 14, 15, 16, 17, 5, 18, 19, 20, 21, 22], and it is now recognized as one of the most viable methods of directly modelling intermediate-mass-ratio inspirals [23] as well as an important tool for calibrating effective models that cover the entire binary parameter space [24, 25, 26].
@@
It is also widely expected that meeting the accuracy requirements of future gravitational-wave detectors [34, 35, 36, 37, 38, 39, 40, 41, 42] will require carrying the SF expansion to second perturbative order in the mass ratio [23, 43], necessitating second-order calculations for generic configurations on a Kerr background. Moreover, at second order in SF theory, the spin of the secondary object enters [44] and contributes significantly to the waveform phase [45]. Over the past several years, this importance of the secondary spin has spurred rapid progress toward SF waveform models for spinning secondaries on generic orbits around spinning primaries [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 45, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. Our aim in this paper is to describe a complete waveform-generation framework for that generic scenario, focusing on the impact of the secondary’s spin.
@@ I.1 Multiscale waveform generation with a nonspinning secondary
If we neglect the secondary’s spin, then generic inspiraling orbits around Kerr black holes are approximately tri-periodic, with three independent, slowly evolving frequencies $\Omega^{i}=(\Omega^{r},\Omega^{\theta},\Omega^{\phi})$ [79]. Each frequency has an associated phase $\mathring{\psi}^{i}=\int\Omega^{i}\,dt$ describing the phase of the secondary’s radial, azimuthal, or polar motion. Calculations are performed in the background Kerr geometry with the primary’s spin axis orthogonal to the equatorial plane $\theta=\pi/2$, such that spin-orbit precession manifests itself as precession of the secondary’s orbital plane (meaning $\Omega^{\theta}\neq\Omega^{r}$).
@@
where $\dot{\cal E}^{(1)}_{\cal H}$ and $\dot{\cal L}^{(1)}_{\cal H}$ are the leading-order fluxes of energy and angular momentum into the primary’s horizon. We use rings over variables that cleanly separate slow evolution from rapid oscillations, avoiding oscillations in the mode amplitudes $\mathring{h}^{(n)}_{\bm{k}}$ and in the right-hand side of Eqs. (2) and (3).
In practical models, all of the inputs, $\mathring{h}^{(n)}_{\bm{k}}$, $F^{(n)}_{i}$, $\dot{\cal E}^{(n)}_{\cal H}$, and $\dot{\cal L}^{(n)}_{\cal H}$, are pre-computed on a grid of $\mathring{\pi}_{i}$ values. [^1] The background spin must also be included as an axis on this parameter-space grid, while the background mass only provides an overall scale. The corrections $\delta m_{1}$ and $\delta\chi_{1}$ do not need to be included as axes because $h^{(2)}_{\bm{k}}$ and $F^{(1)}_{i}$ are linear in these corrections, meaning one can pre-compute coefficients of $\delta m_{1}$ and $\delta\chi_{1}$ as functions of $\mathring{\pi}_{i}$ and of the background spin. Waveforms are then rapidly generated by solving Eqs. (2)–(5) and summing the modes in Eq. (1). Equations (2)–(5) can be solved rapidly because all dependence on the phases $\mathring{\psi}^{i}$ has been eliminated through the choice of variables $(\mathring{\psi}^{i},\mathring{\pi}_{i})$, avoiding the need to resolve oscillations on the orbital time scales $\sim 2\pi/\Omega^{i}$ [81].
In Eqs. (2)–(5), numeric labels in parentheses denote the post-adiabatic (PA) order at which the term enters the evolution of the orbital phases [76, 78], and $F^{(n)}_{i}$ are referred to as nnPA forcing functions. To understand this order counting, note Eqs. (2)–(5) imply that the orbital phases, and hence the waveform phase [22], have an asymptotic expansion of the form
@@
Although this expansion has limited accuracy compared to a direct solution of Eqs. (2)–(5[5], it provides a simple estimate of the impact of any given term in the dynamics: nnPA forcing functions directly contribute to the nnPA phase $\mathring{\psi}_{(n)}^{i}$.
@@ II.1 Field equations and effective metric
$T_{(m)}^{\alpha\beta}$ $=m_{2}\int d\hat{\tau}^{\prime}\,\frac{\delta^{4}\left[x^{\mu}-z^{\mu}(\tau^{\prime})\right]}{\sqrt{-\hat{g}^{\prime}}}\hat{u}^{\alpha}(\tau^{\prime})\hat{u}^{\beta}(\tau^{\prime}),$ (10a)
$T_{(d)}^{\alpha\beta}$ $=(m_{2})^{2}\,\hat{\nabla}_{\rho}\int d\hat{\tau}^{\prime}\,\frac{\delta^{4}\left[x^{\mu}-z^{\mu}(\tau^{\prime})\right]}{\sqrt{-\hat{g}^{\prime}}}\hat{u}^{(\alpha}(\tau^{\prime}){\hat{S}}^{\beta)\rho}(\tau^{\prime}),$ (10b)
@@
Importantly, the stress-energy terms (10a) and (10b) take the form of a spinning particle in a certain effective vacuum metric $\hat{g}_{\alpha\beta}$ rather than in the external background $g_{\alpha\beta}$. The proper time $\hat{\tau}$, four-velocity $\hat{u}^{\alpha}\equiv\frac{dz^{\alpha}}{d\hat{\tau}}$, metric determinant $\hat{g}$, and covariant derivative $\hat{\nabla}_{\alpha}$ are all defined from $\hat{g}_{\alpha\beta}$. The effective metric itself is defined by subtraction of suitably defined, singular self-fields $h^{{\rm S}(n)}_{\alpha\beta}$ from the physical metric. Given the physical metric from Eq. (8), which we can write as
@@ IV.1 Orbital motion
The analysis in Ref. [78] is valid for generic accelerated orbits in Kerr, and we will make extensive use of equations in it. However, our treatment differs from Ref. [78] in that we work with $t$ as our time parameter from the beginning, while Ref. [78] begins with a complete analysis in terms of Mino time $\lambda$ and then transforms to variables based on tt. [^5] We also make several changes of notation from Ref. [78]: $p^{i}\to\pi_{i}$, $\varphi_{i}\to\mathring{\psi}^{i}$, $p^{i}_{\varphi}\to\mathring{\pi}_{i}$, and $\mathscr{P}^{\alpha}\to\varpi_{I}$, along with various less noteworthy alterations. Differential equations in terms of the two variables are related by the factor
@@ IV.1.1 Osculating geodesic equations
$\cos\theta(\psi^{i},\pi_{i})$ $=\sin\iota\cos\psi^{\theta},$ (81b)
@@
$\frac{d\mathring{\psi}^{i}_{(0)}}{dt}=\Omega^{i}_{(0)}(\pi_{i})\qquad\text{(geodesic case)},$ (94)
@@ IV.1.2 Averaging transformation
$\hskip-5.0ptF^{(0)}_{i}+\Omega^{j}_{(0)}\frac{\partial\delta\pi_{i}}{\partial\mathring{\psi}^{j}}+\Omega_{s(0)}\frac{\partial\delta\pi_{i}}{\partial\mathring{\psi}_{s}}$ $=f^{(0)}_{i}(\psi^{j}_{(0)},\mathring{\varpi}_{j},\tilde{\psi}_{s(0)}),$ (97b)
@@
is the average over $\mathring{\psi}^{i}$ and $\mathring{\psi}_{s}$. Note that $\mathring{\psi}^{\phi}$ does not appear in the equations of motion, meaning $\langle\cdot\rangle$ reduces to $\frac{1}{(2\pi)^{3}}\oint\cdot\,d\mathring{\psi}^{r}d\mathring{\psi}^{\theta}d\mathring{\psi}_{s}$ in the above expressions. We also note that the 0PA forcing function $F^{(0)}_{i}$ is independent of $\chi_{2}$, $\delta m_{1}$, and $\delta\chi_{1}$ despite the fact that $f^{(0)}_{i}$ depends on these quantities. This is because the first-order MPD force and the linear force due to $\delta m_{1}$ and $\delta\chi_{1}$ are purely conservative. We remind the reader of this property in more detail in Appendix B.
@@
Note that the criterion of removing oscillations only determines the oscillatory parts of the functions $\Delta\psi^{i}$, $\delta\pi_{i}$, and $\Delta\psi_{s}$ (and their higher-order analogues); their averages $\langle\Delta\psi^{i}\rangle$, $\langle\Delta\pi_{i}\rangle$, and $\langle\Delta\psi_{s}\rangle$ do not enter into equations such as (102). Hence, these non-oscillatory pieces of the transformation can be chosen as arbitrary functions of $\mathring{\pi}_{i}$ (or of $\mathring{\varpi}_{J}$ in the case of $\langle\Delta\pi_{i}\rangle$). However, such a choice affects the subleading terms in the equations of motion (87), (88), and (92). Concretely, we see from Eqs. (103) that the non-oscillatory functions have the following contribution to the subleading frequency corrections and forcing functions:
@@ IV.2 Field equations
$\hskip-5.0pt\left(\frac{\partial}{\partial t}\right)_{\!\bm{k},q}$ $=-i\bigl(k_{i}\Omega^{i}_{(0)}+q\,\Omega_{s(0)}\bigr)$
@@
$T^{(m)}_{\alpha\beta}(\mathring{\psi}^{i},\mathring{\varpi}_{I},\mathring{\psi}_{s},x^{i},\varepsilon)=m_{2}\frac{\hat{g}_{\alpha\mu}\hat{g}_{\beta\nu}\dot{z}^{\mu}\dot{z}^{\nu}}{\sqrt{-\hat{g}_{\rho\sigma}\dot{z}^{\rho}\dot{z}^{\sigma}}}\frac{\delta^{3}(x^{i}-z^{i})}{\sqrt{-\hat{g}}}$ (121)
@@
$\cos\theta_{(0)}$ $=\sin\mathring{\iota}\cos\psi^{\theta}_{(0)},$ (126b)
@@
Given the expansions of the Einstein tensor and stress-energy tensor, the Einstein equation (14) divides into a hierarchical set of first- and second-order equations for the Fourier mode coefficients $\mathring{h}^{(1,\bm{k})}_{\alpha\beta}$ and $\mathring{h}^{(2,\bm{k},q)}_{\alpha\beta}$:
@@
[…]
@@ IV.3 Linear-in-spin effects
where $h^{(2\text{-}\chi_{2})}_{\alpha\beta}$ is the metric perturbation sourced by the spin. The first source term on the right-hand side is confined to the libration region containing the particle’s orbit. The second source term, arising from $\Omega^{i}_{(1\text{-}\chi_{2})}$ and $\Omega_{s(1\text{-}\chi_{2})}$ and Eq. (110), is distributed over the entire spacetime. As highlighted in Ref. [45] and discussed in the next section, there is considerable advantage in choosing a (phase-space) gauge that eliminates this non-compact term.
@@ IV.4.1 Fixed frequencies and fixed turning points
    the phases $\mathring{\psi}^{i}$, which are angular coordinates on a torus in phase space, share an origin with the phases $\psi^{i}$. Concretely, $\mathring{\psi}^{i}(\psi^{j}=0,\mathring{\pi}_{j})=0$, meaning $\mathring{\psi}^{r}=0$ corresponds to periapsis, and $\mathring{\psi}^{\theta}=0$ corresponds to maximum inclination (for orbits with fixed $\mathring{\pi}_{i}$).
@@
    either $\mathring{\pi}_{i}$ is geodesically related to $\Omega_{s}\equiv\frac{d\mathring{\psi}_{s}}{dt}$, making $\Omega_{s}=\Omega_{s(0)}(\mathring{\pi}_{i})$, or $\mathring{\psi}_{s}$ shares an origin with $\tilde{\psi}_{s}$.
We stress that our condition of ‘fixed turning points’ is unrelated to the condition with the same name in Ref. [61]; ours is a condition on the values of the radial and polar phases $\mathring{\psi}^{r}$ and $\mathring{\psi}^{\theta}$ at turning points (periapsis and maximum inclination), while Ref. [61]’s is a condition on the values of Boyer-Lindquist radius $r$ and polar angle $\theta$ at turning points.
@@ V.1 Waveform generation scheme
$h$ $=\sum_{\bm{k}\in\mathbb{Z}^{3}}\Bigl[\varepsilon\mathring{h}^{(1)}_{\bm{k}}(\mathring{\pi}_{i},\theta,\phi)+\varepsilon^{2}\mathring{h}^{(2\text{-}\bcancel{\chi_{2}})}_{\bm{k}}(\mathring{\varpi}_{i},\theta,\phi)$
@@
with $\mathring{h}^{(1)}_{\bm{k}}=\lim_{r\to\infty}\Bigl(rh^{(1,\bm{k})}_{\alpha\beta}\bar{m}^{\alpha}\bar{m}^{\beta}\Bigr)$, for example. Equation (179) extends Eq. (1) to include the secondary spin. Here $\mathring{h}^{(2\text{-}\bcancel{\chi_{2}})}_{\alpha\beta}$ denotes the $\chi_{2}$-independent part of the second-order metric perturbation, identical to the second-order term in Eq. (1). We have also divided the linear spin contribution—the solution to Eq. (152)—into pieces proportional to $\chi_{\parallel}$ and $\chi_{\perp}$, respectively.